OpenStudy (anonymous):

What is the equation of the line that passes through (-3, -1) and has a slope of ? Put your answer in slope-intercept form. y = 3/5x + 4/5 y =3/5 x - 4/5 y = -3/5x -4/5

OpenStudy (anonymous):

ALSO What is the slope of the line that passes through (-2, 7) and (4, 9)? slope = -1/3 slope = 1/3 slope = 3